psirockin2:

Find an equation that models the path of a satellite if its path is a hyperbola. Find an equation that models the path of a satellite if its path is a hyperbola, a = 55,000 km, and c = 81,000 km. Assume that the center of the hyperbola is the origin and the transverse axis is horizontal. I know you're supposed to use the Pythagorean Theorem

psirockin2:

A=(55,000,0) C=(81,000,0) c^2 = a^2 + b^2 81000^2 = 55,000^2 + b^2 I have this so far. I can find B but what do I do next?

psirockin2:

Um... @Hero?

psirockin2:

Anyone?

psirockin2:

Do I just stick A and B into the (x^2/a^2)-(y^2/b^2)=1 formula?

psirockin2:

Nevermind

Nnesha:

yes

psirockin2:

Okay thanks